Circular No. 7683

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2001dp IN NGC 3953 M. Migliardi and E. Dal Farra, Tourtour, Haute-Provence, France,

report their discovery of an apparent supernova (mag 14.5) on CCD images obtained on Aug. 12.9 (limiting mag ∼ 17) and 13.9 UT with a 0.2-m Schmidt-Cassegrain telescope in the course of the CROSS program (cf. IAUC 7373). The new object is located at α = 11^{h}53^{m}45^{s}.1, δ = +52^{o}20′57′′ (equinox 2000.0), which is 24′′ west and 81′′ north of the nucleus of NGC 3953. A CROSS image taken on Apr. 13 showed nothing at this position (limiting mag ∼ 18.5), and it does not appear on Palomar Sky Survey images. Migliardi adds that another image obtained on Aug. 14.83 shows SN 2001dp at mag 14.4. T. Kato, Kyoto University, forwards the following position end figures from an unfiltered CCD image obtained with a 0.28-m reflector by Y. Sano, Nayoro, Hokkaido, Japan, on Aug. 15.491 (supernova at mag 14.6, from USNO-A2.0 red comparison-star magnitudes): 45^{s}.22, 56′′.8 (which is 34′′.8 west and 80′′.2 north of the galaxy nucleus).

K. Ayani, Bisei Astronomical Observatory (BAO), reports that spectra (range 400–800 nm) of SN 2001dp were obtained with the BAO 1.01-m telescope on Aug. 15.5 UT by himself with T. Kawabata, M. Ioroi, O. Ohshima, and participants of the BAO Summer School 2001. The spectra resemble that of the type-Ia SN 1994D, one month after maximum.

SUPERNOVA 2001dq IN IC 1222 M. Modjaz and W. D. Li, University of California at Berkeley, report

the discovery by LOTOSS (cf. IAUC 7514) of an apparent supernova (mag ∼ 18.8) in unfiltered images taken with the Katzman Automatic Imaging Telescope (KAIT) on Aug. 14.2 and 15.2 UT (limiting mag ∼ 20.0). SN 2001dq is also present on earlier images taken on July 28.2, Aug. 1.2, and Aug. 9.2, all with mag ∼ 18.8. The new object is located at α = 16^{h}35^{m}10^{s}.95, δ = +46^{o}12′50′′.5 (equinox 2000.0), which is 18′′.5 east and 0′′.2 north of the nucleus of IC 1222. KAIT images taken on 2001 July 23.2 (limiting mag ∼ 19.0) and 1998 Sept. 13.2 (limiting mag ∼ 19.5) showed nothing at this position.

SUPERNOVA 2001do IN UGC 11459 M. Armstrong, Rolvenden, Kent, England, reports that he obtained

an unfiltered CCD image of SN 2001do (cf. IAUC 7682) at mag 16.2 during his search program on Aug. 14.9 UT. He provides the following precise position: α = 19^{h}37^{m}22^{s}.76, δ = +40^{o}42′22′′.8 (equinox 2000.0)

2001 August 15  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green